OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new system of equations that would result if x was isolated in the first equation. x + 6y + z = 2 x – 3y + 2z = –1 x + 5y + 3z = –4

OpenStudy (mertsj):

x= 2-z-6y 2nd equation becomes: (2-z-6y)-3y+2z = -1 or -9y+z=-3 3rd equation becomes: (2-z-6y)+5y+3z=-4 or -y+2z =-6